# Unlocking Scaling Law in Industrial Recommendation Systems with a Three-step Paradigm based Large User Model

### v1

@@ Abstract.
Recent advancements in autoregressive Large Language Models (LLMs) have achieved significant milestones, largely attributed to their scalability, often referred to as the ”scaling law”. Inspired by these achievements, there has been a growing interest in adapting LLMs for Recommendation Systems (RecSys) by reformulating RecSys tasks into generative problems. However, these End-to-End Generative Recommendation (E2E-GR) methods tend to prioritize idealized goals, often at the expense of the practical advantages offered by traditional Deep Learning based Recommendation Models (DLRMs) in terms of in features, architecture, and practices. This disparity between idealized goals and practical needs introduces several challenges and limitations, locking the scaling law in industrial RecSys. In this paper, we introduce a large user model (LUM) that addresses these limitations through a three-step paradigm, designed to meet the stringent requirements of industrial settings while unlocking the potential for scalable recommendations. Our extensive experimental evaluations demonstrate that LUM outperforms both state-of-the-art DLRMs and E2E-GR approaches. Notably, LUM exhibits excellent scalability, with performance improvements observed as the model scales up to 7 billion parameters. Additionally, we have successfully deployed LUM in an industrial application, where it achieved significant gains in an A/B test, further validating its effectiveness and practicality.
* These authors contributed equally to this work and are co-first authors
$\dagger$ Corresponding author
@@ 1. Introduction
Consequently, several studies have endeavored to mimic the generative models approach by reformulating the retrieval and ranking tasks within RecSys as generative tasks. This is achieved through utilizing User Behavior Sequences (UBS) as the corpus and training a transformer architecture on a ”next-item prediction” task in an end-to-end fashion (Zhai et al., 2024; Geng et al., 2022). However, these end-to-end Generative Recommendation methods (E2E-GRs) tend to overemphasize idealized goals, thereby neglecting the inherent strengths in features, architecture, and practices of DLRMs. This gap between idealized goals and practical implementation in the modeling paradigm has subsequently given rise to a series of challenges and limitations.
(1) Inconsistency Between Generative Training and Discriminative Application: Despite their proficiency in capturing the intricate patterns and distributions inherent in data, E2E-GRs often exhibit limitations when applied to specific discriminative tasks, such as click-through rate prediction where both calibration ability and ranking ability are paramount (Yan et al., 2022). This disparity stems from the intrinsic focus of generative models on the process of data generation rather than on the precise predictive outcomes. As a result, although generative models are adept at capturing inherent distributions of data, they may not consistently satisfy the rigorous demands of discriminative applications, which require high levels of accuracy and specificity (Bernardo et al., 2007; Ng and Jordan, 2001; Yogatama et al., 2017).
(2) Efficiency Challenges: The demand for high efficiency in continuous streaming training, coupled with the stringent low-latency requirements for online real-time inference (beyond the throughput considerations mentioned in (Zhai et al., 2024)), presents significant challenges for the direct implementation of E2E-GRs in industrial settings. Even when deployment is feasible, these efficiency constraints continue to impede the scalability of these E2E models (Section 5.4).
@@
(4) Limited Compatibility: E2E-GRs are constructed using raw UBS with a transformer architecture. This approach inherently limits their compatibility with pre-existing industrial knowledge, such as explicit feature engineering and the parameter inheritance from DLRMs. Consequently, a significant performance gap often emerges between the developmental stages of these E2E-GRs and their real-world deployment in industrial settings (Section 5.3). This gap is particularly pronounced in applications where the online model has been refined over several years or even decades. This discrepancy underscores the need for more robust integration strategies to bridge the gap between theoretical advancements and practical applications.
@@
In summary, our contribution are summarized as follows:
$\bullet$ We introduce a pioneering three-step paradigm specifically tailored for industrial applications.
$\bullet$ We propose a large user model (LUM) that incorporates a ”next-condition-item prediction” task, to bridge the gap between generative pre-training and discriminative applications.
$\bullet$ Comprehensive empirical evaluations demonstrate that the proposed three-step paradigm based LUM significantly enhances the performance of downstream tasks in both public and industrial settings. Notably, our experiments reveal a clear scaling law governing the performance of LUM, which underscores the importance of scaling up the model size for optimal results. Additionally, we have successfully deploy it on our industrial applications and achieve significant gains in an A/B test. This observation not only validates the effectiveness of our approach but also serves as a catalyst for further exploration and advancement in this research direction.
@@ 2. Related works
Deep Learning based Recommendation Models. Traditional DLRMs typically employ a deep neural network and can be broadly categorized into two primary classes: (1) Retrieval-Oriented Models: These models primarily focus on retrieving the most relevant items for users. They often utilize a two-tower architecture, such as EDB (Huang et al., 2020), or sequential models, like SASRec (Kang and McAuley, 2018), BERT4Rec (Sun et al., 2019), and GRU4Rec(Hidasi, 2015) to effectively capture the correlations between users and items. (2) Click-Through Rate (CTR) Prediction Models: This category of models is designed to predict the click-through rate. Many of these models adopt an Embedding + Multi-Layer Perceptron (MLP) architecture. Some works, such as DeepFM (Guo et al., 2017) and xDeepFM (Lian et al., 2018), aim to build more complex interaction signals, while others, including DIN (Zhou et al., 2018), DIEN (Zhou et al., 2019), SIM (Pi et al., 2020) and TWIN (Chang et al., 2023), explore the significance of UBS. Despite the significant improvements achieved by these methods in RecSys, they generally fail to scale with increasing computational resources and do not fully leverage the advancements in large foundation models.
Generative Recommendation Models. To investigate the scaling laws, several studies have been conducted to emulate LLM architectures and construct autoregressive transformers through next-item prediction in an end-to-end manner (Zhai et al., 2024; Geng et al., 2022). However, as discussed in Section 1, these approaches often make strong and idealized assumptions that overlook the advantages of traditional DLRMs in terms of in features, architecture, and practices. This results in a series of issues and limitations when these models are applied to industrial settings. In this paper, we propose a three-step paradigm with LUM to address these limitations. It is worth noting that there are also works that attempt to leverage the open-world knowledge in LLMs to build content-based recommendation models via end-to-end (E2E) or multi-step training (Bao et al., 2023; Lin et al., 2024; Yu et al., 2024; Chen et al., 2024). These efforts, however, fall outside the scope of our current study. Our focus is on developing a scalable model that captures collaborative filtering signals rather than content signals.
@@ 3.2. E2E-GRs via Next-item Prediction
In the context of E2E-GRs, given a UBS for user $u$ represented as $B_{u}=\{i_{1},i_{2},...,i_{L}\}$ where $L$ denotes the length of the behavior sequence, the next-item prediction framework posits that the probability of an item $i_{k}$ appearing next is conditionally dependent on the preceding items $\{i_{1},i_{1},i_{2},...,i_{k-1}\}$. Consequently, the likelihood of the entire UBS $B_{u}$ can be mathematically expressed as: $p(i_{1},i_{2},...,i_{L})=\prod_{l=1}^{L}p(i_{l}|i_{1},i_{2},...,i_{l-1})$. The objective of the autoregressive learning process within E2E-GRs is to optimize the distribution $p_{\theta}(i_{1},i_{2},...,i_{L})$, a method commonly referred to as ”next-item prediction”.
@@
Figure 3. (a) The architecture of LUM. (b) An example of query knowledge from pre-trained LUM. (c) An example of utilizing knowledge in DLRMs.
@@ 4.1.1. Tokenization
In this study, we propose a novel approach to reformulating autoregressive modeling in UBS by transitioning from a traditional ”next-item prediction” framework to a more sophisticated ”next-condition-item prediction” paradigm. Specifically, each item $i$ within the sequence $B_{u}$ is decomposed into two distinct tokens: a condition token and an item token. Consequently, the UBS can be represented as a sequence of alternating tokens: $\{c_{1},i_{1},c_{2},i_{2},...,c_{l},i_{l}\}$, where $c_{k}$ denotes the condition token associated with the item $i_{k}$ (as illustrated in Figure 2). To elucidate, consider a practical scenario where a user’s behavior encompasses interactions with items across multiple scenarios, such as recommendation and search scenarios. In this context, the condition token $c_{i}$ can be defined as a scenario token, thereby capturing the specific environment in which the item $i$ was encountered. This approach facilitates a nuanced understanding of user preferences and behaviors across different scenarios. It is important to note that while other methods, such as HSTU (Zhai et al., 2024), have introduced additional action tokens, their methodology differs fundamentally from our proposed framework. In HSTU, each item is extended as ¡item, action¿ where the action $a_{k}$ is associated with the preceding item $i_{k}$, rather than the subsequent item (Figure 2). As a result, HSTU’s strategy is limited in its ability to predict the next item based on varying conditions and fails to adequately capture user preferences across different aspects.
@@ 4.1.2. Architecture
Token Encoder. The input tokens for LUM exhibit heterogeneity, characterized by two primary categories: condition token $c$ and item token $i$. Moreover, individual tokens may carry diverse attribute features, such as ID, statistical, and content features associated with item tokens. The token encoder is devised to integrate these heterogeneous inputs into a unified token embedding. This process involves initially concatenating the various features or embeddings of each token, followed by a projection layer to consolidate and transform these features into a common representational space. Mathematically, this transformation is formalized as: $e^{t}=\textit{proj}^{t}(\textit{concat}(f^{t}_{1};f^{t}_{2};f^{t}_{3};...));t\in\{i,c\}$ where $f^{t}_{k}$ refers to the features of item token $i$ or condition token $c$, and $e^{t}$ denotes the resultant token embedding. In this study, a linear projection is employed for $proj^{t}$ unless otherwise specified.
User Encoder: The user encoder is structured to capture user preferences and the collaborative information among items. Specifically, the sequence of input tokens $\{c_{1},i_{1},c_{2},i_{2},...,c_{L},i_{L}\}$ is represented as $\{e^{c}_{1},e^{i}_{1},e^{c}_{2},e^{i}_{2},...,e^{c}_{L},e^{i}_{L}\}$ through the token encoder. Subsequently, as illustrated in Figure 3 (a), the user encoder utilizes a conventional autoregressive transformer architecture to process these embeddings. The final output of the user encoder is denoted as $o^{c}_{k}$, encapsulating the integrated information from the input sequence.
@@ 4.1.3. Next-condition-item Prediction
Unlike next-item prediction, which focuses on predicting the subsequent item directly, next-condition-item prediction is concerned with predicting the next item given a specific condition. This approach necessitates the application of an autoregressive loss solely on the output of the condition token to infer the next item. Consequently, the autoregressive likelihood for next-condition-item prediction can be formulated as follows:
(1) $p(c_{1},i_{1},c_{2},i_{2},...,c_{L},i_{L})=\prod_{l=1}^{L}p(i_{l}|c_{1},i_{1},c_{2},i_{2},...,i_{l-1},c_{l})$
Furthermore, to enhance the optimization of $p_{\theta}(c_{1},i_{1},c_{2},i_{2},...,c_{L},i_{L})$ in practical industrial applications, we employ the InfoNCE loss function (Oord et al., 2018) and introduce a packing strategy.
InfoNCE Loss. In industrial applications, the vocabulary size of items can scale to billions, rendering the direct computation of generative probabilities over the entire set of items impractical. To address this challenge, we employ the InfoNCE loss for predicting the next conditional items. The InfoNCE loss can be mathematically formulated as follows:
(2) $Loss=-\sum_{l=1}^{L}\log\left(\frac{\exp(sim(o^{c}_{l-1},e^{i}_{l})))}{\exp(sim(o^{c}_{l-1},e^{i}_{l})+\sum_{k=1}^{K}\exp(sim(o^{c}_{l-1},e^{i}_{k})))}\right)$
where $sim$ is the similarity function. For each item $l$, the other items within the same batch serve as negative samples and $K$ is the number of negative items. $e^{i}_{k}$ is the embedding of $k$-th negative items.
@@ 4.2. Step 2: Knowledge Querying with Given Conditions
In the Step 1, we construct the joint probability distribution denoted as $p(c_{1},i_{1},c_{2},i_{2},...,c_{L},i_{L})$. The next step involves extracting related knowledge from this established probability. Specifically, the tokenization methodology introduced in Section 4.1.1 facilitates the knowledge querying under various conditions. Given a query condition $c_{q}$, the conditional probability $p(i_{q}|c_{1},i_{1},c_{2},i_{2},...,c_{L},i_{L},c_{q})$ can be computed to ascertain the likelihood of a user’s interest in item $i_{q}$ (see Figure 3 (b)). It is noteworthy that this approach to trigger conditional knowledge establishes a bridge between generative models and discriminative tasks, thereby enhancing effectiveness in industrial applications. The following examples illustrate the application of different condition tokens:
Example 1. When the condition token corresponds to a scenario token, the model can infer user interests across different scenarios.
Example 2. If the condition token represents a search query token within search scenarios, the model can deduce user interests based on varying search queries.
Example 3. In cases where the condition token signifies a category token, the model can ascertain user interests in different categories.
Furthermore, the model’s flexibility allows for the simultaneous consideration of multiple types of conditions by incorporating additional condition features into the set $\{f^{c}_{1};f^{c}_{2};f^{c}_{3};...\}$. Empirical evidence demonstrates that integrating diverse conditions significantly enhances performance (see Section 5.3). In essence, this process can be conceptualized as a form of ”prompt engineering” designed to elicit a wide range of knowledge.
@@
Group query for efficiency. Moreover, given that a single user may response multiple queries, each pertaining to the same UBS, processing these queries in isolation can lead to significant inefficiencies. This issue is exacerbated in practical scenarios where the number of users can easily scale to billions, leading to an extensive number of ¡user, query¿ pairs requiring inference. To mitigate this challenge, we introduce a novel group query strategy aimed at enhancing computational efficiency. As illustrated in Figure 4, all queries are concatenated into a single sequence, represented as $p(i_{q_{1}},i_{q_{2}},..|c_{1},i_{1},c_{2},i_{2},...,c_{L},i_{L},c_{q_{1}},c_{q_{2}},...)$. To ensure that the inference process remains coherent and accurate, we apply a masking mechanism to prevent attentional interactions between different query condition $c_{q_{j}}$. This approach allows the common prefix $\{c_{1},i_{1},c_{2},i_{2},...,c_{L},i_{L}\}$ of various queries to be computed only once, while simultaneously querying the items $i_{q_{j}}$ under different conditions. Empirical evaluations demonstrate that, with the implementation of the group query strategy, the inference process can be significantly accelerated by 78%(see Section 5.3).

### v2

@@ Abstract.
Recent advancements in autoregressive Large Language Models (LLMs) have achieved remarkable progress, largely driven by their scalability—commonly formalized as the scaling law. Inspired by these successes, there has been growing interest in adapting LLMs to recommendation systems (RecSys) by reformulating recommendation tasks as generative sequence modeling problems. However, existing End-to-End Generative Recommendation (E2E-GR) methods often sacrifice the practical advantages of traditional Deep Learning-based Recommendation Models (DLRMs)—including mature feature engineering, modular architectures, and production-grade optimization practices. This trade-off introduces critical challenges that hinder the effective application of scaling laws in industrial RecSys. In this paper, we present Large User Model (LUM), a scalable and production-aware framework that bridges the gap between generative modeling and industrial recommendation requirements. LUM addresses these limitations through a principled three-step paradigm, designed to preserve the flexibility of autoregressive generation while maintaining compatibility with real-world deployment constraints. Extensive experiments show that LUM outperforms state-of-the-art DLRMs and E2E-GR approaches across multiple benchmarks. Notably, LUM exhibits strong scalability: performance improves consistently as the model scales up to 7 billion parameters. Furthermore, LUM has been successfully deployed in a large-scale industrial application, where it delivered statistically significant gains in a live A/B test, demonstrating both its effectiveness and practical viability.
Large User Model,Scaling Law,Recommendation Systems
journalyear: 2026copyright: ccconference: Proceedings of the Nineteenth ACM International Conference on Web Search and Data Mining; February 22–26, 2026; Boise, ID, USAbooktitle: Proceedings of the Nineteenth ACM International Conference on Web Search and Data Mining (WSDM ’26), February 22–26, 2026, Boise, ID, USAdoi: 10.1145/3773966.3777970isbn: 979-8-4007-2292-9/2026/02ccs: Information systems Novelty in information retrievalccs: Information systems Sponsored search advertisingccs: Information systems Personalization
@@ 1. Introduction
Consequently, a growing body of work has sought to adapt the generative modeling paradigm to recommendation systems by reformulating retrieval and ranking as generative tasks. This is typically accomplished by treating User Behavior Sequences (UBS) as a behavioral corpus and training a Transformer-based architecture in an end-to-end manner to perform next-item prediction (Zhai et al., 2024; Geng et al., 2022). However, these End-to-End Generative Recommendation methods (E2E-GRs) often disregard the well-established advantages of traditional Deep Learning-based Recommendation Models (DLRMs)—including rich feature engineering, modular architectures, and production-proven training practices. This design choice introduces a critical trade-off, leading to a range of practical challenges that limit their effectiveness and scalability in industrial settings.
(1) Inconsistency Between Generative Training and Discriminative Application: Despite their ability to capture complex sequential patterns and underlying data distributions, E2E-GRs often underperform on discriminative downstream tasks—such as click-through rate (CTR) prediction—where both calibration and ranking performance are critical (Yan et al., 2022a). This performance gap arises from a fundamental misalignment: generative models are optimized to maximize data likelihood, focusing on accurately modeling the process of user behavior generation, rather than directly optimizing for discriminative objectives like prediction accuracy or ranking quality. Consequently, while E2E-GRs excel at distributional modeling, they may fail to meet the stringent requirements of industrial recommendation pipelines, which demand highly calibrated probabilities and fine-grained ranking sensitivity (Ng and Jordan, 2001; Bernardo et al., 2007; Yogatama et al., 2017).
(2) Efficiency Challenges: The demand for high efficiency in continuous streaming training, coupled with stringent low-latency requirements for online real-time inference—extending beyond the throughput considerations discussed in (Zhai et al., 2024)—presents significant challenges for the direct deployment of E2E-GRs in industrial settings. Even when deployment is feasible, these efficiency constraints fundamentally limit model scalability, particularly in terms of both model capacity and input sequence length (see Section 5.4).
@@
(4) Limited Compatibility. E2E-GRs are typically built on raw UBS using a Transformer architecture. This design inherently limits their compatibility with established industrial practices—such as explicit feature engineering and parameter inheritance from deployed DLRMs—hindering seamless integration into production systems. As a result, a significant performance gap often arises between the development phase of E2E-GRs and their real-world deployment (Section 5.3). This gap is particularly evident in applications where the online model has been iteratively refined over years or even decades. Such a discrepancy highlights the need for frameworks that better bridge the divide between theoretical advances and industrial practice.
@@
In summary, our contribution are summarized as follows: (1) We introduce a pioneering three-step paradigm specifically tailored for industrial applications. (2) We propose a large user model (LUM) that incorporates a ”next-condition-item prediction” task, to bridge the gap between generative pre-training and discriminative applications. (3) Comprehensive empirical evaluations demonstrate that the proposed three-step paradigm based LUM significantly enhances the performance of downstream tasks in both public and industrial settings. Notably, our experiments reveal a clear scaling law governing the performance of LUM, which underscores the importance of scaling up the model size for optimal results. Additionally, we have successfully deploy it on our industrial applications and achieve significant gains in an A/B test.
@@ 2. Related works
Deep Learning based Recommendation Models. Traditional DLRMs typically use deep neural networks and fall into two categories: (1) Retrieval-oriented models, such as two-tower architectures (e.g., EDB (Huang et al., 2020)) or sequential models (e.g., SASRec (Kang and McAuley, 2018), BERT4Rec (Sun et al., 2019), GRU4Rec (Hidasi, 2015)), which capture user-item relevance; (2) CTR prediction models, often based on Embedding + MLP structures (Guo et al., 2017; Lian et al., 2018; Zhou et al., 2018, 2019; Pi et al., 2020; Chang et al., 2023; Yan et al., 2022b). Despite their success, these models struggle to scale with compute and fail to harness advances in large foundation models.
Generative Recommendation Models. To study scaling laws, prior work has adapted LLM-like autoregressive transformers for next-item prediction (Zhai et al., 2024; Geng et al., 2022). However, as noted in Section 1, these methods rely on idealized assumptions and neglect the strengths of traditional DLRMs in features, architecture, and industrial practices—limiting their real-world applicability. We address this with a three-step paradigm based on LUM. While some recent efforts leverage LLMs’ open-world knowledge for content-based recommendations via end-to-end or multi-step training (Bao et al., 2023; Lin et al., 2024; Yu et al., 2024; Chen et al., 2024), our focus remains on scalable modeling of collaborative signals, not content.
@@ 3.2. E2E-GRs via Next-item Prediction
In the context of E2E-GRs, given a UBS for user $u$, denoted as $B_{u}=\{i_{1},i_{2},\dots,i_{L}\}$, where $L$ represents the length of the sequence, the next-item prediction framework assumes that the probability of the next item $i_{k}$ is conditionally dependent on the preceding items $\{i_{1},i_{2},\dots,i_{k-1}\}$. Accordingly, the likelihood of the entire UBS $B_{u}$ can be factorized as: $p(i_{1},i_{2},\dots,i_{L})=\prod_{l=1}^{L}p(i_{l}\mid i_{1},i_{2},\dots,i_{l-1}).$ The objective of the autoregressive learning paradigm in E2E-GRs is to model and optimize the joint distribution $p_{\theta}(i_{1},i_{2},\dots,i_{L})$, a task commonly referred to as ”next-item prediction” [^1] It is worth noting that some approaches extend this formulation by jointly modeling the types of user actions (e.g., click, purchase, view) (Zhai et al., 2024). However, the underlying modeling principle remains fundamentally similar..
@@
Figure 3. The three-step paradigm based large user model.
@@ 4.1.1. Tokenization
In this work, we propose a novel reformulation of autoregressive modeling in UBS, shifting from the conventional ”next-item prediction” framework to a more expressive ”next-condition-item prediction” paradigm. Specifically, each item $i_{k}$ in the sequence $B_{u}=\{i_{1},\dots,i_{L}\}$ is decomposed into two distinct tokens: a condition token $c_{k}$ and an item token $i_{k}$. As a result, the UBS is transformed into an alternating sequence: $\{c_{1},i_{1},c_{2},i_{2},\dots,c_{L},i_{L}\},$ where $c_{k}$ encodes contextual signals associated with the subsequent item $i_{k}$ (see Figure 2). To illustrate, consider a user who interacts with items across multiple application scenarios—such as search and recommendation. In this case, $c_{k}$ can be instantiated as a scenario token (e.g., ”search” or ”homepage recommendation”), explicitly modeling the context in which $i_{k}$ was engaged. This design addresses a key limitation of standard autoregressive models: without explicit conditioning, generated items reflect a general interest distribution that may be misaligned with downstream tasks. For instance, considering a search scenario, after a user issues the query ”dresses”, a model lacking condition control may generate recommendations unrelated to fashion or intent. In contrast, our framework enables explicit control over generation behavior: during inference, the condition token acts as a trigger that steers the model toward task-specific outputs by setting appropriate context. This mechanism is conceptually analogous to prompt engineering in large language models, where different prompts elicit distinct knowledge patterns.
Notably, while HSTU (Zhai et al., 2024) also uses auxiliary tokens (e.g., action types), its modeling differs fundamentally: each interaction is a tuple $\langle i_{k},a_{k}\rangle$, where action $a_{k}$ (e.g., click) is tied to the past item $i_{k}$ (Figure 2), not used to condition future predictions. This limits its ability to dynamically shape the next-item distribution based on prospective context. In contrast, our formulation supports context-aware generation and task-controllable inference, significantly enhancing E2E-GR flexibility in practice.
@@ 4.1.2. Architecture
Token Encoder. The input tokens in LUM are inherently heterogeneous, falling into two primary types: condition tokens ($c$) and item tokens ($i$). Each token type is associated with a set of attribute features—for instance, item tokens may include ID-based (e.g., item ID, brand ID), statistical (e.g., click-through rate), and content-derived (e.g., title embedding) features, while condition tokens may incorporate contextual metadata such as scenario ID or query text. To unify these heterogeneous inputs, we design a token encoder that maps each token into a shared embedding space. The encoding process first concatenates the feature representations of a token, followed by a projection layer that transforms the concatenated vector into a compact, dense embedding. Formally, the embedding $e^{t}$ of token type $t\in\{i,c\}$ is computed as: $e^{t}=\textit{proj}^{t}(\textit{concat}(f^{t}_{1};f^{t}_{2};f^{t}_{3};\dots));t\in\{i,c\}$ where $f^{t}_{k}$ denotes the $k$-th feature representation of token $t$, and $\textit{proj}^{t}$ is a learnable transformation. In this work, $\textit{proj}^{t}$ is implemented as a linear layer unless otherwise specified.
User Encoder: The user encoder is structured to capture user preferences and the collaborative information among items. Specifically, the sequence of input tokens $\{c_{1},i_{1},c_{2},i_{2},\dots,c_{L},i_{L}\}$ is represented as $\{e^{c}_{1},e^{i}_{1},e^{c}_{2},e^{i}_{2},\dots,e^{c}_{L},e^{i}_{L}\}$ through the token encoder. Subsequently, as illustrated in Figure 3 (a), the user encoder utilizes a conventional autoregressive transformer architecture to process these embeddings. The final output of the user encoder is denoted as $o^{c}_{k}$, encapsulating the integrated information from the input sequence.
@@ 4.1.3. Next-condition-item Prediction
Unlike next-item prediction, which focuses on predicting the subsequent item directly, next-condition-item prediction is concerned with predicting the next item given a specific condition. This approach necessitates the application of an autoregressive loss solely on the output of the condition token to infer the next item. Consequently, the autoregressive likelihood for next-condition-item prediction can be formulated as follows: $p(c_{1},i_{1},c_{2},i_{2},\dots,c_{L},i_{L})=\prod_{l=1}^{L}p(i_{l}|c_{1},i_{1},c_{2},i_{2},\dots,i_{l-1},c_{l})$ Furthermore, to enhance the optimization of $p_{\theta}(c_{1},i_{1},c_{2},i_{2},\dots,c_{L},i_{L})$ in practical industrial applications, we employ the InfoNCE loss function (Oord et al., 2018) and introduce a packing strategy.
InfoNCE Loss. In industrial applications, the vocabulary size of items can scale to billions, rendering the direct computation of generative probabilities over the entire set of items impractical. To address this challenge, we employ the InfoNCE loss for predicting the next conditional items. The InfoNCE loss can be mathematically formulated as follows: $Loss=-\sum_{l=1}^{L}\log\left(\frac{\exp(sim(o^{c}_{l},e^{i}_{l})))}{\exp(sim(o^{c}_{l},e^{i}_{l})+\sum_{k=1}^{K}\exp(sim(o^{c}_{l},e^{i}_{k})))}\right)$ where simsim is the similarity function. For each item $l$, the other items within the same batch serve as negative samples and $K$ is the number of negative items. $e^{i}_{k}$ is the embedding of $k$-th negative items.
@@ 4.2. Step 2: Knowledge Querying with Given Conditions
In Step 1, we construct the joint probability distribution over the tokenized sequence: $p(c_{1},i_{1},c_{2},i_{2},\dots,c_{L},i_{L}),$ as introduced in Section 4.1.1. The subsequent step involves querying task-relevant knowledge from this distribution through conditional generation. Specifically, for each downstream task request, we first construct a corresponding condition token $c_{q}$ based on the query context. For example, if the request corresponds to a user searching for items related to the query “dresses” in a search scenario, we can set $c_{q}$ to include the associated scenario ID and query text. Given $c_{q}$, the conditional probability $p(i_{q}\mid c_{1},i_{1},c_{2},i_{2},\dots,c_{L},i_{L},c_{q})$ is computed to estimate the likelihood of user interest in item $i_{q}$ (see Figure 3(b)).
Notably, this mechanism—using condition tokens as triggers to elicit specific knowledge—establishes a bridge between generative sequence modeling and discriminative downstream tasks, thereby improving adaptability and effectiveness in real-world applications. We illustrate this capability through the following examples:
$\bullet$ Example 1. When $c_{q}$ encodes a scenario identifier (e.g., “search” or “homepage”), the model infers user interests under different application contexts.
$\bullet$ Example 2. If $c_{q}$ represents a search query, the model generates recommendations conditioned on fine-grained user intent.
$\bullet$ Example 3. When $c_{q}$ denotes a category label, the model captures category-specific preferences.
Moreover, the framework supports multi-condition input by incorporating multiple feature fields into the condition token (i.e., extending the set $\{f^{c}_{1};f^{c}_{2};f^{c}_{3};\dots\}$), enabling joint conditioning on scenario, intent, and metadata. Empirical results show that integrating diverse conditions significantly improves performance (see Section 5.3). Conceptually, this process mirrors prompt engineering in large language models, where carefully designed prompts activate relevant knowledge patterns—here, condition tokens serve as structured prompts to steer the generative model toward task-specific outputs.
It is noteworthy that, in Step 2, the condition token may encounter cold-start issues when dealing with new scenarios or new categories, which cannot be directly triggered. As for new queries, since we obtain embeddings from tokenized query text as features for the condition token, the textual features inherently possess a certain level of generalization to alleviate the cold-start problem. Considering the relatively low frequency of new scenarios and categories in industrial applications, along with the flexibility to adjust the contextual content of the condition token, the cold-start issue is not the focus of this study and will be left for future work.
@@
Group query for efficiency. Moreover, given that a single user may response multiple queries, each pertaining to the same UBS, processing these queries in isolation can lead to significant inefficiencies. This issue is exacerbated in practical scenarios where the number of users can easily scale to billions, leading to an extensive number of ¡user, query¿ pairs requiring inference. To mitigate this challenge, we introduce a novel group query strategy aimed at enhancing computational efficiency. As illustrated in Figure 4, all queries are concatenated into a single sequence, represented as $p(i_{q_{1}},i_{q_{2}},\dots|c_{1},i_{1},c_{2},i_{2},\dots,c_{L},i_{L},c_{q_{1}},c_{q_{2}},\dots)$. To ensure that the inference process remains coherent and accurate, we apply a masking mechanism to prevent attentional interactions between different query condition $c_{q_{j}}$. This approach allows the common prefix $\{c_{1},i_{1},c_{2},i_{2},\dots,c_{L},i_{L}\}$ of various queries to be computed only once, while simultaneously querying the items $i_{q_{j}}$ under different conditions. Empirical evaluations demonstrate that, with the implementation of the group query strategy, the inference process can be significantly accelerated by 78%(see Section 5.3).